OpenStudy (anonymous):

There are two fruit trees located at (3,0) and (–3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.

OpenStudy (anonymous):

medal and fan!!!

OpenStudy (anonymous):

@IMStuck hey can you help

OpenStudy (anonymous):

@zepdrix hi can you help

OpenStudy (anonymous):

\((3,0)\) and \((–3, 0)\) are to be the focal points of the ellipse right?

OpenStudy (anonymous):

for this you want \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] and you want \[a^2-b^2=3^2\] the easiest way to do that is to use the famous \(3-4-5\) right triangle and make \[a=5,b=4\] so \(c=3\) and use \[\frac{x^2}{5^2}+\frac{y^2}{4^2}=1\] that will make your foci \((-3,0)\) and \((3,0)\)

OpenStudy (anonymous):

check the nice picture here to see that it works http://www.wolframalpha.com/input/?i=ellipse+x^2%2F25%2By^2%2F16%3D1

OpenStudy (anonymous):

for the hyperbola it will look like \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] and you want \(a^2+b^2=3^3\) simplest way i can think to do it is to make \(a^2=8,b^2=1\) and use \[\frac{x^2}{8}-y^2=1\] but you have other choices

OpenStudy (anonymous):

Perfect!! thank you @satellite73

OpenStudy (anonymous):

here is a nice graph with both together if you need one http://www.wolframalpha.com/input/?i=+x^2%2F8-y^2%3D1%2Cx^2%2F25%2By^2%2F16%3D1

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

why is the formula a^2+b^2=3^3 ? @satellite73

OpenStudy (anonymous):

why is a very good question it will take a lot of explaining however a better explanation is here http://www.purplemath.com/modules/hyperbola2.htm

OpenStudy (anonymous):

well not why... how

OpenStudy (anonymous):

can you please explain the hyperbola part

OpenStudy (anonymous):

@hero can you explain the hyperbola part please